# Can We Hide Machines in the Crowd? Quantifying Equivalence in LLM-in-the-loop Annotation Tasks

### v1

@@ 4.2. Data Selection
For evaluation, we selected 38 coders split into two groups of 19 coders each: We used the elbow point finding method detailed in Section 3.3 to determine the optimal number of annotators in a group. As shown in Figure 4, this point was at 19 annotators per group. We divided the coders into two groups:
@@ 4.3. LLM Substitution Evaluation
To assess variability and statistical significance, we applied the paired bootstrap procedure described in Section 3.1.4. We tested across multiple bootstrap sample sizes, with $B\in\{50,100,\ldots,500\}$. For each value of $B$, we repeated the substitution experiment 10 times. Each repetition involved generating $B$ bootstrap samples of 40 items per iteration, computing agreement scores, and performing the Two One-Sided t-Test (TOST) as described in Section 3.2.
For each value of $B$, we report the mean and standard deviation over the 10 trials for the two p-values from the TOST procedure ($p_{1}$ and $p_{2}$).
@@ 5. Discussion
In the PolitiFact dataset, replacing human ratings with LLM ratings lowers reliability across raters (see Figure 10). Fact-checking is harder than movie rating because it needs domain knowledge, careful use of evidence, and awareness of political framing. Human annotators bring in their own beliefs, expertise, and even mood, which creates variability (Roitero et al., 2020; Jensen et al., 2016; Gilbert et al., 1993). Different cues like familiarity, political views, or emotional reactions influence people in different ways (Hammond, 1955; Hoffman, 1960). The LLM, however, takes a narrower and cautious approach: it avoids strong labels like “True” or “False” and stays near the middle categories Figure 8. This mismatch with human judgment patterns explains why inter-rater reliability drops more in this case.
The alpha-change analysis (Figures 9 and 10) shows a clear contrast. In MovieLens, swapping some human raters with the LLM barely affects reliability, meaning the model can mimic certain annotators well. In PolitiFact, though, replacements cause big drops in agreement for some raters. This suggests LLMs still have trouble replicating the unique, knowledge-based reasoning humans use in complex or disputed topics.
@@
Practitioners should apply our evaluation methodology as follows. For existing datasets, replicate the workflow described in Section 4. If the dataset does not meet our minimum criteria (four annotators, at least 80 items), a small-scale annotation effort should be undertaken. The same approach applies to new tasks: collect a modest but diverse sample of annotations, compute reliability measures, and test LLM substitution feasibility before scaling.
@@ 6. Conclusion
This method provides a practical way to detect differences in annotation behavior between humans and LLMs. It also offers an opportunity to evaluate a small set of annotations before deciding whether to use an LLM for large-scale annotation in a specific application.
@@
The performance of the LLM is likely to vary depending on the specific model and the prompts used. Future work should explore how different LLM architectures and prompt strategies influence the results, enabling a more robust evaluation of LLM-human alignment across annotation contexts.

### v2

@@ 4.2. Data Selection
For evaluation, we selected 38 coders split into two groups of 19 coders each. We used the elbow point finding method detailed in Section 3.3 to determine the optimal number of annotators in a group. As shown in Figure 5, this point was at 19 annotators per group. We divided the coders into two groups:
@@ 4.3. LLM Substitution Evaluation
To assess variability and statistical significance, we applied the paired bootstrap procedure described in Section 3.1.4. We tested multiple bootstrap sample sizes, with $B\in\{50,100,\ldots,500\}$. For each value of $B$, we repeated the substitution experiment 10 times. Each repetition involved generating $B$ bootstrap samples of 40 items per iteration, computing agreement scores, and performing the Two One-Sided t-Test (TOST) as described in Section 3.2. For each value of $B$, we report the mean and standard deviation over the 10 trials for the two p-values from the TOST procedure ($p_{1}$ and $p_{2}$).
@@ 5. Discussion
In the PolitiFact dataset, replacing human ratings with LLM ratings lowers reliability across raters (see Figure 11). Fact-checking is harder than movie rating because it requires domain knowledge, evidence usage, and political framing awareness. Human annotators bring in their own beliefs, expertise, and even mood, which creates variability (Roitero et al., 2020; Jensen et al., 2016; Gilbert et al., 1993). Different cues like familiarity, political views, or emotional reactions influence people in different ways (Hammond, 1955; Hoffman, 1960). The LLM, however, takes a narrower and cautious approach: it avoids strong labels like “True” or “False” and stays near the middle categories Figure 9. This mismatch with human judgment patterns explains why inter-rater reliability drops more in this case.
The alpha-change analysis (Figures 10 and 11) shows a clear contrast. In MovieLens, swapping some human raters with the LLM has a minimal impact on reliability, indicating that the model can effectively mimic certain annotators. In PolitiFact, though, replacements cause big drops in agreement for some raters. This suggests LLMs still have trouble replicating the unique, knowledge-based reasoning humans use in complex or disputed topics.
@@
Practitioners should apply our evaluation methodology as follows. For existing datasets, replicate the workflow described in Section 4. If the dataset does not meet our minimum criteria (four annotators, at least 80 items), a small-scale annotation effort should be undertaken. The same approach applies to new tasks: collect modest and diverse annotations, compute reliability measures, and test LLM substitution feasibility before scaling.
@@ 6. Conclusion
This method provides a practical way to detect differences in annotation behavior between humans and LLMs. It also offers an opportunity to evaluate a small set of annotations before deciding whether to use an LLM for large-scale annotations.
@@
The performance of the LLM is likely to vary depending on the specific model and the prompts used. Future work should investigate how various LLM architectures and prompt strategies impact the results, facilitating a more comprehensive evaluation of LLM-human alignment across diverse annotation contexts.